Please help me solve this: A rectangle is 4 times as long as it is wide. If the length is increased by 4 inches and the width is decreased by 1 inch,the area will be 60 square inches. What were the dimensions of the original rectangle?
Answer by mananth(16946)

let width be w
length = 4w
length increased by 4=>4w+4
width decreased by 1=>w-1
Area = 60
Area = l*w
(w-1)(4w+4)=60
(w+1)*4(w-1)=60
(w+1)(w-1)=60/4
w^2-1=15
w^2=15+1
w^2=16
w=+/-4
ignore negative length
w=4 in
length = 4*4 = 16in